# MIA-Bench: Towards Better Instruction Following Evaluation of Multimodal LLMs

### v3

@@ 3.2 Supervised Fine-Tuning (SFT)
LLaVA-NeXT-13b [30] 69.16 86.75 69.88 82.07 64.77 74.99 48.56 77.50 75.83
LLaVA-NeXT-13b* 78.85 86.90 86.80 88.02 71.34 81.01 60.87 84.17 72.65
Table 3: Detailed results on MIA-Bench before and after (denoted by *) supervised fine-tuning on additional constructed diverse instruction-tuning data. We re-ran the baseline.
Model MMBench TextVQA VQA2 LLaVA-itw POPE VizWiz MathVista MIA-Bench
LLaVA-NeXT-13b [30] 70.6 64.26 82.80 85.8 87.7 60.41 33.0 69.16
LLaVA-NeXT-13b* 68.6 63.20 82.58 83.4 86.9 59.72 32.0 78.85
Table 4: Results on MIA-Bench and other major multimodal benchmarks before and after (denoted by *) supervised fine-tuning on additional diverse instruction-tuning data. We re-ran the baseline.
@@
Results. Using LLaVA-NeXT-13b as the backbone, we train the model for 1 epoch on the constructed SFT data. Results on MIA-Bench and other benchmarks are summarized in Table 4, with detailed results on MIA-Bench reported in Table 3. The performance on MIA-Bench has been significantly improved by around 10 points, at the cost of minor regressions across other benchmarks. Examples are shown in Figure 7 to compare responses from LLaVA-NeXT-13b before and after SFT.
@@ 4 Related Work
Multimodal LLMs and Benchmarks. Multimodal Large Language Models (MLLMs) have recently emerged as a significant research focus. LLaVA [4] and MiniGPT-4 [47] pioneered visual instruction tuning, and the past year has witnessed a boom of open-source MLLMs based on this concept. Prominent examples include InstructBLIP [5], mPLUG-Owl(-2/Doc) [40, 48, 49], Qwen-VL [7], CogVLM [8], SPHINX(-X) [9, 50], InternLM-XComposer2-VL [51], InternVL(-1.5) [52, 2], VILA [53], MM1 [11], Mini-Gemini [54], Idefics2 [55], Phi-3-vision [31], to name a few. There is also a rich body of literature on enabling MLLMs for referring and grounding [56, 57, 58, 59, 60, 61, 62, 63], image generation and editing [64, 65, 66], etc.
@@
Various benchmarks have been proposed to evaluate the performance of MLLMs across different dimensions. Benchmarks like VQAv2 [12], TextVQA [34], ScienceQA [14], MME [15], MMbench [45], SEED-Bench [17], MathVista [18], and MMMU [3] aim to assess comprehensive multimodal understanding abilities. Additionally, there are benchmarks that specifically study model hallucination, including POPE [67], MHalDetect [68], GAVIE [69], HallusionBench [46], and MAD-Bench [70]. Many of these benchmarks have gained popularity within the community due to their use of multiple-choice evaluations. However, they do not accurately reflect the common use cases for MLLMs, where user interactions are typically open-ended. To address this, benchmarks like LLaVA-Bench [4], MM-Vet [19], and Vibe-Eval [22] have been proposed. Our MIA-Bench also falls into this category; however, we focus on studying the exact instruction adherence of MLLMs, a metric that previous benchmarks have only loosely measured.
@@
Instruction Following Benchmarks for LLMs. Several benchmarks have been proposed to measure the instruction adherence ability of LLMs. Instruction-Following Eval (IFEval) [25] is a benchmark for assessing LLMs’ adherence ability to the given instructions. Its approach emphasizes verifiable instructions, which enhance objectivity and reproducibility in evaluations. IFEval creates 541 prompts spanning 25 instruction types, revealing a significant performance gap in instruction adherence ability between GPT-4 [1] and PaLM-2 [71]. This demonstrates the benchmark’s ability to effectively differentiate between models in adherence ability. On the other hand, InfoBench [26] introduces a new metric called Decomposed Requirements Following Ratio (DRFR) for assessing the instruction-adherence capabilities of LLMs. DRFR dissects complex instructions into simpler sub-instructions, allowing for a granular evaluation of compliance with various task aspects. InfoBench contains 500 diverse instructions consisting of 2,250 decomposed questions in multiple constraint categories. The evaluation of advanced LLMs using this framework highlights their strengths and areas for improvement, especially in complex instruction adherence scenarios. Compared with these previous work, we are the first known effort that specifically focuses on benchmarking the instruction adherence ability of multimodal LLMs.
@@ 5 Conclusion
This paper introduces MIA-Bench, a benchmark designed to evaluate the ability of MLLMs to strictly adhere to complex instructions within prompts. Through the analysis of 400 image-prompt pairs from diverse sources, our findings highlight variability in model performance and much room for improvement, underscoring a critical need for enhanced training methods to improve instruction compliance. We further explored supervised fine-tuning (SFT) using LLaVA-NeXT as the backbone, which yielded promising results. Going forward, future research can expand on both SFT and alignment methods such as RLHF [72, 73] and DPO [74, 75], enhancing MLLMs to achieve higher accuracy and reliability in practical applications across diverse instructional contexts.
@@ Appendix A Appendix
A.1 Comparison between MIA-Bench and IFEval
Figure 8: Comparison between MIA-Bench and IFEval.
A.2 Examples of diverse instruction-tuning data used for SFT
@@
A.3 Examples of how MLLMs respond to instructions in MIA-Bench

### v5

@@ 3.2 Supervised Fine-Tuning (SFT)
LLaVA-NeXT-13b (Liu et al., 2024a) 69.16 86.75 69.88 82.07 64.77 74.99 48.56 77.50 75.83
      + DIT 78.85 86.90 86.80 88.02 71.34 81.01 60.87 84.17 72.65
      + DIT + LLaVA-Instruct150k 78.90 88.59 74.67 79.95 74.17 66.39 53.70 100.00 80.83
Table 4: Detailed results of LLaVA-NeXT-13b (Liu et al., 2024a) on MIA-Bench before and after supervised fine-tuning on additional constructed diverse instruction-tuning (DIT) data, and the mixture of DIT and LLaVA-Instruct150k. We re-ran the baseline.
Model MMBench TextVQA VQA2 LLaVA-itw POPE VizWiz MIA-Bench
LLaVA-NeXT-13b (Liu et al., 2024a) 70.6 64.26 82.80 85.8 87.7 60.41 69.16
      + DIT 68.6 63.20 82.58 83.4 86.9 59.72 78.85
      + DIT + LLaVA-Instruct150k 67.27 54.24 77.92 75.8 87.8 58.87 78.90
Table 5: Results of LLaVA-NeXT-13b (Liu et al., 2024a) on MIA-Bench and other major multimodal benchmarks supervised fine-tuning on additional constructed diverse instruction-tuning (DIT) data, and the mixture of DIT and LLaVA-Instruct150k. We re-ran the baseline.
@@
Results. Using LLaVA-NeXT-13b as the backbone, we train the model for 1 epoch on the constructed diverse instruction-tuning (DIT) data. We also performed SFT using the combination of LLaVA Visual Instruct 150K dataset and our diverse instruction-tuning dataset, to examine which data mixture leads to better performance. Results on MIA-Bench and other benchmarks are summarized in Table 5, with detailed results on MIA-Bench reported in Table 4. The performance of the model after finetuning on the DIT data alone on MIA-Bench has been significantly improved by around 10 points, at the cost of minor regressions across other benchmarks. This supports our hypothesis that using the small amount of DIT data, we can enhance the model with better instruction following capability. On the other hand, further adding LLaVA Visual Instruct 150K data did not significantly improve model performance on the tested benchmarks. Examples are shown in Figure 8 to compare responses from LLaVA-NeXT-13b before and after SFT.
@@ 4 Related Work
Multimodal LLMs and Benchmarks. Multimodal Large Language Models (MLLMs) have recently emerged as a significant research focus. LLaVA (Liu et al., 2023c) and MiniGPT-4 (Zhu et al., 2023) pioneered visual instruction tuning, and the past year has witnessed a boom of open-source MLLMs based on this concept. Prominent examples include InstructBLIP (Dai et al., 2023), mPLUG-Owl(-2/Doc) (Ye et al., 2023b; c; a), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023a), SPHINX(-X) (Lin et al., 2023b; Gao et al., 2024), InternLM-XComposer2-VL (Dong et al., 2024b), InternVL(-1.5) (Chen et al., 2023c; 2024b), VILA (Lin et al., 2023a), MM1 (McKinzie et al., 2024), Mini-Gemini (Li et al., 2024b), Idefics2 (Laurençon et al., 2024), Phi-3-vision (Abdin et al., 2024), to name a few. There is also a rich body of literature on enabling MLLMs for referring and grounding (Peng et al., 2023b; Chen et al., 2023a; You et al., 2024a; Wang et al., 2023b; Lai et al., 2023; Zhang et al., 2023; 2024; You et al., 2024b), image generation and editing (Koh et al., 2023; Sun et al., 2023a; Fu et al., 2023b), etc.
@@
Various benchmarks have been proposed to evaluate the performance of MLLMs across different dimensions. Benchmarks like VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019a), ScienceQA (Lu et al., 2022), MME (Fu et al., 2023a), MMbench (Liu et al., 2024b), SEED-Bench (Li et al., 2023a), MathVista (Lu et al., 2023), and MMMU (Yue et al., 2023) aim to assess comprehensive multimodal understanding abilities. Additionally, there are benchmarks that specifically study model hallucination, including POPE (Li et al., 2023b), MHalDetect (Gunjal et al., 2024), GAVIE (Liu et al., 2023a), HallusionBench (Guan et al., 2023), and MAD-Bench (Qian et al., 2024). Many of these benchmarks have gained popularity within the community due to their use of multiple-choice evaluations. However, they do not accurately reflect the common use cases for MLLMs, where user interactions are typically open-ended. To address this, benchmarks like LLaVA-Bench (Liu et al., 2023c), MM-Vet (Yu et al., 2023b), and Vibe-Eval (Padlewski et al., 2024) have been proposed. Our MIA-Bench also falls into this category; however, we focus on studying the exact instruction adherence of MLLMs, a metric that previous benchmarks have only loosely measured.
@@
Instruction Following Benchmarks for LLMs. Several benchmarks have been proposed to measure the instruction adherence ability of LLMs. Instruction-Following Eval (IFEval) (Zhou et al., 2023) is a benchmark for assessing LLMs’ adherence ability to the given instructions. Its approach emphasizes verifiable instructions, which enhance objectivity and reproducibility in evaluations. IFEval creates 541 prompts spanning 25 instruction types, revealing a significant performance gap in instruction adherence ability between GPT-4 (Achiam et al., 2023) and PaLM-2 (Anil et al., 2023). This demonstrates the benchmark’s ability to effectively differentiate between models in adherence ability. On the other hand, InfoBench (Qin et al., 2024) introduces a new metric called Decomposed Requirements Following Ratio (DRFR) for assessing the instruction-adherence capabilities of LLMs. DRFR dissects complex instructions into simpler sub-instructions, allowing for a granular evaluation of compliance with various task aspects. InfoBench contains 500 diverse instructions consisting of 2,250 decomposed questions in multiple constraint categories. The evaluation of advanced LLMs using this framework highlights their strengths and areas for improvement, especially in complex instruction adherence scenarios. Compared with these previous work, we are the first known effort that specifically focuses on benchmarking the instruction adherence ability of multimodal LLMs.
@@ 5 Conclusion
This paper introduces MIA-Bench, a benchmark designed to evaluate the ability of MLLMs to strictly adhere to complex instructions within prompts. Through the analysis of 400 image-prompt pairs from diverse sources, our findings highlight variability in model performance and much room for improvement, underscoring a critical need for enhanced training methods to improve instruction compliance. We further explored supervised fine-tuning (SFT) using LLaVA-NeXT as the backbone, which yielded promising results. Going forward, future research can expand on both SFT and alignment methods such as RLHF (Ouyang et al., 2022; Sun et al., 2023b) and DPO (Rafailov et al., 2024; Yu et al., 2023a), enhancing MLLMs to achieve higher accuracy and reliability in practical applications across diverse instructional contexts.
@@
Reproducibility Statement
For reproducibility purpose, we release our evaluation code and benchmark at: https://github.com/apple/ml-mia-bench.
@@ Appendix A Appendix
A.1 Examples of diverse instruction-tuning data used for SFT
@@
A.2 Examples of how MLLMs respond to instructions in MIA-Bench
@@
A.3 Comparison of scores and rankings across different judge models
GPT-4o 0.813587 0.823910 0.847312 0.965594 0.733660 0.820370 0.775862 0.733333 0.765351
Clause-3-Opus 0.789474 0.805156 0.794086 0.916667 0.697115 0.771364 0.706897 0.629630 0.692308
Reka 0.776965 0.805536 0.807151 0.857904 0.732143 0.740253 0.687500 0.600000 0.689189
MiniCPM-Llama3-V2.5 0.737190 0.715403 0.799068 0.849505 0.735294 0.711585 0.637931 0.716667 0.693694
Gemini-1.0-Pro 0.674504 0.676852 0.764286 0.768350 0.697695 0.615801 0.683333 0.611111 0.588235
LLaVA-1.5-13b 0.615427 0.663542 0.643424 0.643873 0.423077 0.619357 0.359195 0.666667 0.449561
ShareGPT4v 0.602989 0.674028 0.601496 0.625556 0.538462 0.560248 0.456897 0.555556 0.547619
Idefics-2-8b 0.442778 0.437893 0.494687 0.461218 0.520408 0.436042 0.276786 0.458333 0.412162
Table 6: Details of model scores evaluated by gpt-4o-mini-2024-07-18.
GPT-4o 0.909704 0.927875 0.912371 0.942057 0.862434 0.900441 0.857143 0.916667 0.882883
Clause-3-Opus 0.856077 0.890721 0.868490 0.917070 0.774590 0.819386 0.861111 0.725000 0.809524
Reka 0.839905 0.894752 0.785088 0.905643 0.713661 0.801667 0.925926 0.657407 0.828571
MiniCPM-Llama3-V2.5 0.798023 0.828916 0.771795 0.823087 0.751944 0.763976 0.721264 0.816667 0.841880
Gemini-1.0-Pro 0.773569 0.817422 0.735470 0.788911 0.797814 0.683020 0.866071 0.870370 0.806373
LLaVA-1.5-7b 0.683947 0.758817 0.703750 0.674046 0.630208 0.617620 0.425287 0.800000 0.602564
ShareGPT4v 0.689046 0.800461 0.657738 0.608733 0.654762 0.601754 0.500000 0.800000 0.743056
Idefics-2-8b 0.541755 0.560243 0.619318 0.489276 0.646825 0.455342 0.405556 0.375000 0.627193
Table 7: Details of model scores evaluated by gpt-4o-2024-05-13.
GPT-4o 0.899410 0.909379 0.916204 0.969395 0.854885 0.861247 0.920290 0.907407 0.878378
Clause-3-Opus 0.848949 0.861543 0.871686 0.896552 0.797170 0.808777 0.846154 0.645833 0.865741
Reka 0.826844 0.881841 0.809259 0.873276 0.725309 0.770225 0.814103 0.750000 0.819444
MiniCPM-Llama3-V2.5 0.787537 0.818813 0.790246 0.795796 0.768182 0.736359 0.676667 0.716667 0.828125
Gemini-1.0-Pro 0.763240 0.814379 0.750000 0.785159 0.757682 0.672255 0.758333 0.785714 0.776042
LLaVA-1.5-7b 0.660472 0.751873 0.661822 0.649851 0.498512 0.572719 0.516667 0.750000 0.571429
ShareGPT4v 0.657186 0.765309 0.632682 0.575578 0.583333 0.545104 0.464286 0.675000 0.717742
Idefics-2-8b 0.536134 0.589964 0.541887 0.455882 0.611582 0.449821 0.406667 0.527778 0.576190
Table 8: Details of model scores evaluated by gpt-4o-2024-11-20.
GPT-4o 0.896893 0.906288 0.917996 0.955952 0.830508 0.867949 0.846667 0.833333 0.896396
Claude-3-Opus 0.861628 0.895363 0.866039 0.927730 0.807018 0.820549 0.857639 0.666667 0.800926
Reka 0.830885 0.869867 0.821685 0.883403 0.795597 0.772894 0.813218 0.675000 0.848485
MiniCPM-Llama3-V2.5 0.780966 0.831197 0.766026 0.796257 0.726190 0.722037 0.691358 0.656250 0.768018
Gemini-1.0-Pro 0.757733 0.793860 0.724138 0.745455 0.854167 0.670349 0.810606 0.822917 0.822581
LLaVA-1.5-7b 0.667826 0.743137 0.638889 0.675287 0.571212 0.594505 0.500000 0.758333 0.596774
ShareGPT4v 0.666092 0.773905 0.661290 0.573904 0.562500 0.570722 0.458333 0.638889 0.695238
Idefics-2-8b 0.535057 0.597963 0.531810 0.483768 0.593056 0.452361 0.326087 0.458333 0.569444
Table 9: Details of model scores evaluated by chatgpt-4o-latest.